OpenStudy (anonymous):

If the measure of an exterior angle of a regular polygon is 120°, how many sides does the polygon have?

jimthompson5910 (jim_thompson5910):

Measure of exterior angle = 360/number of sides 120 = 360/n 120n = 360 n = 360/120 n = 3 So there are 3 sides

OpenStudy (anonymous):

ok, thanks you very much!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

hang on, would the formula be 360/n?

jimthompson5910 (jim_thompson5910):

yes, if you have an n-gon, then the measure of one exterior angle is 360/n

OpenStudy (anonymous):

ok ,thanks.